OpenStudy (anonymous):

Simplify. 10.5 • (–3) ÷ 10 • (–2) ÷ 0.5 A. 12.6 B. 3.15 C. –3.15 D. –12.6

OpenStudy (anonymous):

@SingleSixxs_Fiance

OpenStudy (anonymous):

@CamoJunkieQT

OpenStudy (anonymous):

A. 12.6

OpenStudy (anonymous):

C-13.5

OpenStudy (anonymous):

?

OpenStudy (anonymous):

wat is it

OpenStudy (anonymous):

I am correct 10.5/(-3)=-31.5 -31.5/10=-3.15 -3.15*-2= 6.3 6.3/0.5=12.6 Tah-Dah!!!!

OpenStudy (anonymous):

ok thx